41) Teal Mountain Inc. earns $1308000 and pays cash dividends for $392400 during 2017. Bramble Corporation owns 74250 of the 225000 outstanding shares of Teal Mountain. How much revenue from investment should Bramble report in 2017?

$131640.

$431640.

$526560.

$263280.
Answer #1

Share outstanding = 74250/225000 = 33%

Revenue from investment = 1308000*33% = 431640

So answer is b) $431640